OpenStudy (anonymous):

Robert is 15 years older than his brother Peter. However, y years ago Robert was twice as old as Peter. If Peter is now b years old and b > y, find the value of b - y.

OpenStudy (zpupster):

P+15=R 2P =R when we subtract the first equation from the second 2P =R -P+ - 15= -R P-15=0 add 15 both sides P=15 at y years Robert was 30 and Peter was 15 which fits here i am not sure what the question is asking Peter is b meaning older than when he was at y which was 15 so anything older than 15 fits b-y i guess anything >0